OpenStudy (anonymous):

10,000,000,000 = 1010 1011 111 110

OpenStudy (anonymous):

is there a context to this question?

OpenStudy (anonymous):

seems unlikely that ten billion equals one hundred and ten for example

OpenStudy (anonymous):

perhaps the answers have exponents? like \[10^{10}\] or \[10^{11}\]

OpenStudy (zehanz):

There probably should be a ^-sign somewhere in the answers...

OpenStudy (anonymous):

its 10,000,000,000 = 10^10 10^11 1^11 1^10 that better

OpenStudy (zehanz):

You probably know that: \(10=10^1\) \(100=10\cdot10=10^2\) \(1,000=10\cdot10\cdot10=10^3\) etc. Now do some counting yourself with the number 10,000,000,000!